Solve for E and g simultaneously in the following equations:

40×4g×40^3=bw^3E.............(1)

60×4g×60^3=bw^3E.............(2)


1
Expert's answer
2021-08-11T09:53:39-0400

There is no solution:


"\frac{40\cdot4g\cdot40^3}{60\cdot4g\cdot60^3}=\frac{32}{81}\\
\frac{bw^3E}{bw^3E}=1"